OpenStudy (anonymous):

CAL 2Evaluate by using integration by parts

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\sqrt{x} \tan^{-1} (\sqrt{x})dx\]

OpenStudy (anonymous):

\[u=\tan^{-1} (\sqrt{x}) , du=1/1+x*(1/2\sqrt{x})\]

OpenStudy (anonymous):

\[t=\sqrt{x} so dt=1/2\sqrt{x} than 2dt=1/\sqrt{x}\]

OpenStudy (anonymous):

du=1/(x^1/2*(2x+2))